DIRC Operations Report

Wednesday 2007/12/19 to Wednesday 2007/12/26


Oncall: Rolf Andreassen (Jose Benitez on backup)
Fast Monitoring: Rolf Andreassen, Jose Benitez

Contents:

    PEP Activities     BaBar Activities     DIRC activities     Backgrounds     Scheduled activities     Pending Work     Links

Luminosities this week

PEP peak lumi: 5.5 x 1033 (last week: 3.5 x 1033)

PEP delivery: 1.84 fb-1 (last week: 0.57 fb-1)

BaBar integrated lumi:1.88 fb-1 (last week: 0.53 fb-1)

Run 7 total delivery: 2.41 fb-1

Run 7 total BaBar recorded: 2.41 fb-1 total, including 0 fb-1 off-peak data

PEP activities/issues this week

  • PEP mainly tried to push up the lumi; some accesses were neecssary, eg installing an RGD in PR04 on the 20th.
  • On Friday the 21st, the beam energy was changed to run at the Upsilon(3S) resonance.

    BaBar activities/issues since last DIRC meeting

  • Some triggers were changed to better deal with the Upsilon(3S) running; no doubt this will be tuned further as we get more experience.

    DIRC activities/issues this week

  • The chiller alarms continued in spite of our shorting out the relay, and we decided to mask them again in ALH.

    See DIRC logbook for details

    Lumi and Background this week